what is the probability of rolling a 4, 5 or 6 on one fair six sided die?
Answer by ewatrrr(24785)

 
Hi,
what is the probability of rolling a 4, 5 or 6 on one fair 'six sided' die?
P(rolling a 4, 5 'or' 6)= 1/6 + 1/6 + 1/6 = 3/6 = 1/2